The general expression for consecutive multiples of 5 is 5N, 5(N+1), 5(N+2), and so on, where N is some unspecified integer. Find four consecutive multiples of 5 ssuch that 6 times the first is 40 times greater than 2 times the sum of the second and fourth.
Answer by stanbon(75887)

The general expression for consecutive multiples of 5 is 5N, 5(N+1), 5(N+2), and so on, where N is some unspecified integer. Find four consecutive multiples of 5 ssuch that 6 times the first is 40 greater than 2 times the sum of the second and fourth.
---------------------------
EQUATION:
6*5n = 2[(5(n+1)+5(n+3)]+40
30n = 2[10n+20]+40
30n = 20n+80
10n = -80
n = -8
5n=-40
5(n+1)=-35
5(n+2)=-30
5(n+3)=-25
================
Comment: I changed your problem from "40 times greater than" to
40 greater than.
===================
Cheers,
Stan H.